# ANLS* - A Universal Document Processing Metric for Generative Large Language Models

### v9

@@ 4.2 Experimental Evaluation
Dataset Method gpt-4-o gpt-4.5 gemini-1.5-pro mistral-large claude-37 Llama-3.1-405B
DocVQA Simple 0.72 0.760 0.655 0.445 0.763 0.653
Latin Prompting 0.742 0.794 0.734 0.447 0.823 0.595
SFT (Ours) 0.809 0.855 0.752 0.648 0.834 0.737
MPDocVQA Simple 0.623 0.677 0.643 0.364 0.713 0.625
Latin Prompting 0.736 0.767 0.755 0.335 0.780 0.640
SFT (Ours) 0.735 0.777 0.794 0.476 0.790 0.753
Kleister Charity Simple 0.723 0.765 0.748 0.652 0.778 0.665
Latin Prompting 0.736 0.775 0.747 0.576 0.778 0.643
SFT (Ours) 0.752 0.775 0.764 0.657 0.786 0.680
Kleister NDA Simple 0.626 0.638 0.613 0.637 0.629 0.608
Latin Prompting 0.635 0.637 0.608 0.624 0.616 0.591
SFT (Ours) 0.625 0.641 0.623 0.641 0.64 0.610
SROIE Simple 0.905 0.855 0.686 0.855 0.935 0.853
Latin Prompting 0.912 0.836 0.672 0.863 0.877 0.779
SFT (Ours) 0.897 0.877 0.681 0.905 0.899 0.879
VRDU AD Buy Simple 0.645 0.668 0.545 0.386 0.625 0.453
Latin Prompting 0.721 0.725 0.579 0.435 0.639 0.546
SFT (Ours) 0.768 0.726 0.734 0.594 0.641 0.700
VRDU Registration Simple 0.628 0.664 0.549 0.579 0.658 0.614
Latin Prompting 0.628 0.667 0.555 0.587 0.669 0.632
SFT (Ours) 0.624 0.666 0.585 0.639 0.673 0.614
@@ Appendix A Results GPT
Dataset Method gpt-3.5-turbo-16k gpt-3.5-turbo gpt-4 gpt-4-turbo gpt-4-vision gpt-4o gpt-4o + vision gpt-4o-mini gpt-4o gpt-4.5
Version 0613 0125 1106-preview 2024-04-09 preview-1106 2024-05-13 2024-05-13 2024-07-18 2024-08-06 preview-2025-02-27
DocVQA Simple 0.576 0.550 0.607 0.589 0.759 0.656 - 0.568 0.72 0.760
Latin Prompting 0.620 0.626 0.699 0.631 - 0.718 - 0.623 0.742 0.794
SFT (Ours) 0.768 0.733 0.790 0.770 - 0.747 0.754 0.754 0.809 0.855
MPDocVQA Simple 0.507 0.461 0.635 0.606 0.708 0.686 - 0.577 0.623 0.677
Latin Prompting 0.489 0.507 0.739 0.656 - 0.718 - 0.584 0.736 0.767
SFT (Ours) 0.724 0.731 0.781 0.760 - 0.762 0.789 0.731 0.735 0.777
Kleister Charity Simple 0.527 0.564 0.743 0.708 0.751 0.711 - 0.647 0.723 0.765
Latin Prompting 0.477 0.493 0.735 0.728 - 0.721 - 0.672 0.736 0.775
SFT (Ours) 0.534 0.548 0.763 0.733 - 0.737 0.749 0.707 0.752 0.775
Kleister NDA Simple 0.361 0.575 0.695 0.633 0.664 0.649 - 0.622 0.626 0.638
Latin Prompting 0.412 0.593 0.705 0.643 - 0.646 - 0.633 0.635 0.637
SFT (Ours) 0.315 0.562 0.703 0.659 - 0.655 0.648 0.628 0.625 0.641
SROIE Simple 0.509 0.706 0.835 0.849 0.834 0.887 - 0.863 0.905 0.855
Latin Prompting 0.723 0.708 0.851 0.847 - 0.863 - 0.836 0.912 0.836
SFT (Ours) 0.792 0.649 0.873 0.890 - 0.897 0.892 0.872 0.897 0.877
VRDU AD Buy Simple 0.414 0.084 0.553 0.483 0.640 0.596 - 0.398 0.645 0.668
Latin Prompting 0.424 0.211 0.586 0.506 - 0.655 - 0.436 0.721 0.725
SFT (Ours) 0.676 0.452 0.770 0.712 - 0.753 0.755 0.693 0.768 0.726
VRDU Registration Simple 0.586 0.615 0.676 0.609 0.665 0.605 - 0.594 0.628 0.664
Latin Prompting 0.606 0.630 0.673 0.614 - 0.611 - 0.608 0.628 0.667
SFT (Ours) 0.637 0.654 0.711 0.590 - 0.605 0.615 0.596 0.624 0.666
@@ Appendix B Results Gemini
Dataset Method gemini-1.0-pro gemini-1.5-pro-preview gemini-1.5-pro gemini-1.5-pro + vision gemini-1.5-flash gemini-2.0-flash
Version 001 0409 001 001 001 001
DocVQA Simple 0.586 0.647 0.655 - 0.64 0.689
Latin Prompting 0.676 0.668 0.734 - 0.707 0.773
SFT (Ours) 0.741 0.737 0.752 0.850 0.821 0.770
MPDocVQA Simple 0.603 0.627 0.643 - 0.696 0.680
Latin Prompting 0.502 0.726 0.755 - 0.665 0.680
SFT (Ours) 0.616 0.794 0.794 0.870 0.786 0.816
Kleister Charity Simple 0.583 0.764 0.748 - 0.720 0.735
Latin Prompting 0.478 0.763 0.747 - 0.732 0.753
SFT (Ours) 0.633 0.779 0.764 0.695 0.752 0.754
Kleister NDA Simple 0.623 0.722 0.613 - 0.592 0.622
Latin Prompting 0.599 0.707 0.608 - 0.596 0.623
SFT (Ours) 0.552 0.715 0.623 0.520 0.600 0.620
SROIE Simple 0.263 0.594 0.686 - 0.689 0.737
Latin Prompting 0.371 0.623 0.672 - 0.664 0.658
SFT (Ours) 0.288 0.623 0.681 0.931 0.699 0.683
VRDU AD Buy Simple 0.510 0.645 0.545 - 0.597 0.674
Latin Prompting 0.556 0.686 0.579 - 0.628 0.706
SFT (Ours) 0.685 0.765 0.734 0.545 0.701 0.746
VRDU Registration Simple 0.699 0.730 0.549 - 0.593 0.650
Latin Prompting 0.740 0.749 0.555 - 0.633 0.620
SFT (Ours) 0.720 0.780 0.585 0.570 0.667 0.613

### v10

@@ 4.2 Experimental Evaluation
Dataset Method gpt-4.1 gpt-4.5 gemini-2.5-pro mistral-large claude-37 Llama-3.1-405B
DocVQA Simple 0.678 0.760 0.755 0.445 0.763 0.653
Latin Prompting 0.727 0.794 0.803 0.447 0.823 0.595
SFT (Ours) 0.746 0.855 0.863 0.648 0.834 0.737
MPDocVQA Simple 0.671 0.677 0.744 0.364 0.713 0.625
Latin Prompting 0.738 0.767 0.826 0.335 0.780 0.640
SFT (Ours) 0.762 0.777 0.841 0.476 0.790 0.753
Kleister Charity Simple 0.776 0.765 0.755 0.652 0.778 0.665
Latin Prompting 0.773 0.775 0.767 0.576 0.778 0.643
SFT (Ours) 0.764 0.775 0.765 0.657 0.786 0.680
Kleister NDA Simple 0.602 0.638 0.594 0.637 0.629 0.608
Latin Prompting 0.606 0.637 0.605 0.624 0.616 0.591
SFT (Ours) 0.607 0.641 0.610 0.641 0.64 0.610
SROIE Simple 0.906 0.855 0.696 0.855 0.935 0.853
Latin Prompting 0.914 0.836 0.705 0.863 0.877 0.779
SFT (Ours) 0.900 0.877 0.717 0.905 0.899 0.879
VRDU AD Buy Simple 0.672 0.668 0.708 0.386 0.625 0.453
Latin Prompting 0.725 0.725 0.742 0.435 0.639 0.546
SFT (Ours) 0.754 0.726 0.737 0.594 0.641 0.700
VRDU Registration Simple 0.703 0.664 0.597 0.579 0.658 0.614
Latin Prompting 0.710 0.667 0.601 0.587 0.669 0.632
SFT (Ours) 0.664 0.666 0.593 0.639 0.673 0.614
@@ Appendix A Results GPT
Dataset Method gpt-3.5-turbo-16k gpt-3.5-turbo gpt-4 gpt-4-turbo gpt-4-vision gpt-4o gpt-4o + vision gpt-4o-mini gpt-4o gpt-4.5 gpt-4.1 gpt-4.1-mini gpt-4.1-nano
Version 0613 0125 1106-preview 2024-04-09 preview-1106 2024-05-13 2024-05-13 2024-07-18 2024-08-06 preview-2025-02-27 2025-04-14 2025-04-14 2025-04-14
DocVQA Simple 0.576 0.550 0.607 0.589 0.759 0.656 - 0.568 0.72 0.760 0.678 0.616 0.640
Latin Prompting 0.620 0.626 0.699 0.631 - 0.718 - 0.623 0.742 0.794 0.727 0.672 0.713
SFT (Ours) 0.768 0.733 0.790 0.770 - 0.747 0.754 0.754 0.809 0.855 0.746 0.750 0.760
MPDocVQA Simple 0.507 0.461 0.635 0.606 0.708 0.686 - 0.577 0.623 0.677 0.671 0.584 0.632
Latin Prompting 0.489 0.507 0.739 0.656 - 0.718 - 0.584 0.736 0.767 0.738 0.629 0.707
SFT (Ours) 0.724 0.731 0.781 0.760 - 0.762 0.789 0.731 0.735 0.777 0.762 0.691 0.764
Kleister Charity Simple 0.527 0.564 0.743 0.708 0.751 0.711 - 0.647 0.723 0.765 0.776 0.755 0.649
Latin Prompting 0.477 0.493 0.735 0.728 - 0.721 - 0.672 0.736 0.775 0.773 0.770 0.639
SFT (Ours) 0.534 0.548 0.763 0.733 - 0.737 0.749 0.707 0.752 0.775 0.764 0.765 0.683
Kleister NDA Simple 0.361 0.575 0.695 0.633 0.664 0.649 - 0.622 0.626 0.638 0.602 0.623 0.595
Latin Prompting 0.412 0.593 0.705 0.643 - 0.646 - 0.633 0.635 0.637 0.606 0.626 0.579
SFT (Ours) 0.315 0.562 0.703 0.659 - 0.655 0.648 0.628 0.625 0.641 0.607 0.624 0.586
SROIE Simple 0.509 0.706 0.835 0.849 0.834 0.887 - 0.863 0.905 0.855 0.906 0.856 0.804
Latin Prompting 0.723 0.708 0.851 0.847 - 0.863 - 0.836 0.912 0.836 0.914 0.867 0.817
SFT (Ours) 0.792 0.649 0.873 0.890 - 0.897 0.892 0.872 0.897 0.877 0.900 0.879 0.867
VRDU AD Buy Simple 0.414 0.084 0.553 0.483 0.640 0.596 - 0.398 0.645 0.668 0.672 0.651 0.548
Latin Prompting 0.424 0.211 0.586 0.506 - 0.655 - 0.436 0.721 0.725 0.725 0.675 0.581
SFT (Ours) 0.676 0.452 0.770 0.712 - 0.753 0.755 0.693 0.768 0.726 0.754 0.748 0.647
VRDU Registration Simple 0.586 0.615 0.676 0.609 0.665 0.605 - 0.594 0.628 0.664 0.703 0.681 0.562
Latin Prompting 0.606 0.630 0.673 0.614 - 0.611 - 0.608 0.628 0.667 0.710 0.682 0.583
SFT (Ours) 0.637 0.654 0.711 0.590 - 0.605 0.615 0.596 0.624 0.666 0.664 0.646 0.587
@@ Appendix B Results Gemini
Dataset Method gemini-1.0-pro gemini-1.5-pro-preview gemini-1.5-pro gemini-1.5-pro + vision gemini-1.5-flash gemini-2.0-flash gemini-2.5-pro gemini-2.5-flash
Version 001 0409 001 001 001 001 preview-03-25 preview-04-17
DocVQA Simple 0.586 0.647 0.655 - 0.64 0.689 0.755 0.750
Latin Prompting 0.676 0.668 0.734 - 0.707 0.773 0.803 0.826
SFT (Ours) 0.741 0.737 0.752 0.850 0.821 0.770 0.863 0.814
MPDocVQA Simple 0.603 0.627 0.643 - 0.696 0.680 0.744 0.684
Latin Prompting 0.502 0.726 0.755 - 0.665 0.680 0.826 0.809
SFT (Ours) 0.616 0.794 0.794 0.870 0.786 0.816 0.841 0.817
Kleister Charity Simple 0.583 0.764 0.748 - 0.720 0.735 0.755 0.759
Latin Prompting 0.478 0.763 0.747 - 0.732 0.753 0.767 0.768
SFT (Ours) 0.633 0.779 0.764 0.695 0.752 0.754 0.765 0.775
Kleister NDA Simple 0.623 0.722 0.613 - 0.592 0.622 0.594 0.621
Latin Prompting 0.599 0.707 0.608 - 0.596 0.623 0.605 0.605
SFT (Ours) 0.552 0.715 0.623 0.520 0.600 0.620 0.610 0.612
SROIE Simple 0.263 0.594 0.686 - 0.689 0.737 0.696 0.696
Latin Prompting 0.371 0.623 0.672 - 0.664 0.658 0.705 0.671
SFT (Ours) 0.288 0.623 0.681 0.931 0.699 0.683 0.717 0.684
VRDU AD Buy Simple 0.510 0.645 0.545 - 0.597 0.674 0.708 0.651
Latin Prompting 0.556 0.686 0.579 - 0.628 0.706 0.742 0.682
SFT (Ours) 0.685 0.765 0.734 0.545 0.701 0.746 0.737 0.717
VRDU Registration Simple 0.699 0.730 0.549 - 0.593 0.650 0.597 0.620
Latin Prompting 0.740 0.749 0.555 - 0.633 0.620 0.601 0.592
SFT (Ours) 0.720 0.780 0.585 0.570 0.667 0.613 0.593 0.577